cindy has $20,000 to invest. A part of this money is invested in bonds paying 10% annual interest. The rest of the money is invested in a savings account giving 5% annual interest. At the end of the year, she received $1,600 as an extra income. How much money did Cindy place in each investment?
Answer by mananth(16946)

Part I 10.00% per annum ------------- Amount invested =x
Part II 5.00% per annum ------------ Amount invested = y
20000
Interest----- 1600.00

Part I 10.00% per annum ---x
Part II 5.00% per annum ---y
Total investment
x + 1 y= 20000 -------------1
Interest on both investments
10.00% x + 5.00% y= 1600
Multiply by 100
10 x + 5 y= 160000.00 --------2
Multiply (1) by -10
we get
-10 x -10 y= -200000.00
Add this to (2)
0 x -5 y= -40000
divide by -5
y = 8000
Part I 10.00% $ 12000
Part II 5.00% $ 8000

CHECK
12000 --------- 10.00% ------- 1200.00
8000 ------------- 5.00% ------- 400.00
Total -------------------- 1600.00

m.ananth@hotmail.ca